OpenStudy (juan1857):

Find an equation of the tangent line to the graph of f at the given point:

OpenStudy (juan1857):

\[f(x)=x+\frac{ 4 }{ x }, (4,5)\]

OpenStudy (juan1857):

I'm just having trouble finding the derivative

OpenStudy (kropot72):

I suggest you write the equation as: \[f(x)=x+4x^{-1}\] \[\large f'(x)=1-\frac{4}{x^{2}}\]

OpenStudy (juan1857):

can you explain why?

OpenStudy (kropot72):

When you rewrite the equation as above, you can then differentiate by using the usual rule for differentiation.

OpenStudy (juan1857):

is there a formula for that?

OpenStudy (kropot72):

The rule for differentiation of powers of x is: \[\large f(x)=x^{n}\rightarrow\ f'(x)=nx^{n-1}\]

OpenStudy (kropot72):

Have you studied Calculus?

OpenStudy (juan1857):

I am currently studying the course

OpenStudy (juan1857):

can you explain how the denominator became part of the numerator

OpenStudy (kropot72):

Do you understand that: \[\large \frac{1}{x}=x^{-1}\]

OpenStudy (juan1857):

yeah

OpenStudy (juan1857):

how did the x^2 get on bottom

jimthompson5910 (jim_thompson5910):

Power Rule \[\Large f(x) = x^n\] \[\Large f \ '(x) = nx^{n-1}\] where n is a constant

jimthompson5910 (jim_thompson5910):

In this case, n = -1 \[\Large f(x) = x^{-1}\] \[\Large f \ '(x) = -1x^{-1-1}\] \[\Large f \ '(x) = -1x^{-2}\] \[\Large f \ '(x) = -\frac{1}{x^{2}}\]

OpenStudy (kevin):

Then you find gradient : m = f'(4) \[m = 1 - \frac{ 4 }{ x^2 }\] Then, you just need to input (x,y) (y - y1) = m (x -x1) (y - 5) = m (x - 4) Done..........

OpenStudy (canimcan):

Determine the equation of surface which is formed by the tangents of ellipsoid x^2 + y^2/2+z^2=1. Also that surface contain point M(5,5,5). Help me with this. It is the same problem.

OpenStudy (juan1857):

sorry I can't @canimcan

OpenStudy (juan1857):

still need help

OpenStudy (loser66):

with what?

OpenStudy (juan1857):

i don't understand where x^2 came from and how it became the denominator

OpenStudy (loser66):

ok, go steps f(x) = x+4/x hey, you need stay to get help. If you leave, I leave too.

OpenStudy (loser66):

Now, derivative of f(x) =?

OpenStudy (juan1857):

1 +?

OpenStudy (loser66):

ok, tell me 4/x =??

OpenStudy (loser66):

\(\dfrac{4}{x}=4x^{-1}\) ok?

OpenStudy (juan1857):

yeah i understand up to that point

OpenStudy (loser66):

Now, take derivative of it. Apply the n rule.

OpenStudy (loser66):